Determine the electrical force of attraction between two balloons with separate charges of +4.5 x 10^-8 C and -1.9 x 10^-8 C when separated a distance of 0.95 m.




1
Expert's answer
2022-02-28T10:39:58-0500

Given:

"q_1=4.5*10^{-8}\:\rm C"

"q_2=-1.9*10^{-8}\:\rm C"

"r=0.95\:\rm m"


The Coulomb's law says



"F=k\frac{|q_1||q_2|}{r^2}""F=9*10^9*\frac{4.5*10^{-8}*1.9*10^{-8}}{0.95^2}=8.53*10^{-6} \:\rm N"